OpenStudy (anonymous):

What is the supplement of angle (theta symbol) = (pi) / 9 ?

OpenStudy (campbell_st):

well supplementary angles add to 180 degrees, since you are working in radians the would need to add to pi.... so\[\frac{\pi}{9} + ? = \pi\]